OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (-2, -3) and has a slope of 1/2

OpenStudy (anonymous):

1) 2x + y = 4 2) 2x - y = 4 3) x + 2y = 4 4) x - 2y = 4

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ -3}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{1}{2} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and leave 4 on the right}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

hoq do i do that? @jdoe0001